# The single edge notch fracture test for viscoelastic elastomers

### v1

@@ Abstract
Rubber; Elastomers; Adhesives; Cavitation; Fracture
@@ 1 Introduction
As a direct consequence of two elementary observations, Shrimali and Lopez-PamiesShrimali and Lopez-Pamies (2023a) have shown that the original formRivlin and Thomas (1953); Greensmith and Thomas (1955)
$-\dfrac{\partial\mathcal{W}}{\partial\mathrm{\Gamma}_{0}}=T_{c}$ (1)
@@
$-\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial\mathrm{\Gamma}_{0}}=G_{c}.$ (2)
In expression (1), the left-hand side $-\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ denotes the change [^2] The derivative $-\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ is to be carried out under conditions of fixed deformation of those parts of the boundary that are not traction-free so that no work is done by the external forces. in total deformation — stored [^3] The stored part of the energy is comprised of an equilibrium part $\mathcal{W}^{{\rm Eq}}$ and a non-equilibrium part $\mathcal{W}^{{\rm NEq}}$, the latter representing the part of the stored energy that gets dissipated via viscous deformation eventually. and dissipated — energy
$\mathcal{W}=\underbrace{\mathcal{W}^{{\rm Eq}}+\mathcal{W}^{{\rm NEq}}}_{\text{stored}}+\underbrace{\mathcal{W}^{v}}_{\text{dissipated}}$
in the specimen at hand with respect to an added surface area ${\rm d}\mathrm{\Gamma}_{0}$ to the pre-existing crack $\mathrm{\Gamma}_{0}$ in its reference state, while the right-hand side $T_{c}$ stands for the critical tearing energy, a characteristic property of the elastomer that depends on the loading history. Traditionally, the critical tearing energy has been written in the form Gent and Schultz (1972); Gent (1996)
$T_{c}=G_{c}(1+f_{c}),$
@@
In expression (2), on the other hand, the left-hand side $-\partial\mathcal{W}^{{\rm Eq}}/$ $\partial\mathrm{\Gamma}_{0}$ denotes the change not in $\mathcal{W}$ but only in equilibrium elastic energy $\mathcal{W}^{{\rm Eq}}$ stored in the specimen with respect to an added surface area to the pre-existing crack $\mathrm{\Gamma}_{0}$, while the right-hand side involves not $T_{c}$ but only the intrinsic fracture energy $G_{c}$ of the elastomer.
@@
where the right-hand side is to be evaluated at the instance in time at which (2) is attained along the loading path of interest. This last relation clarifies how $T_{c}$ describes the total energy (per unit undeformed fracture area) expended in the tearing process, as it makes explicit the contributions by the creation of new surface ($G_{c}$) and by viscous effects ($-\partial\mathcal{W}^{{\rm NEq}}/\partial\mathrm{\Gamma}_{0}-\partial\mathcal{W}^{v}/\partial\mathrm{\Gamma}_{0}$).
@@
Figure 1: Schematic of the single edge notch fracture test for a viscoelastic elastomer illustrating the geometry of the specimen, the applied time-dependent separation $l(t)$ between the grips, and the corresponding force $P(t)$ at the grips. It is standard practice to use the normalized quantities $\mathrm{\Lambda}=l(t)/L$ and $S=P(t)/(HB)$. We refer to them, respectively, as the global stretch and the global stress. Importantly, the initial size $A$ of the crack (the “notch”) can be equal to, larger, or smaller than the initial thickness $B$ of the specimen.
As a first application of the criticality condition (2), Shrimali and Lopez-PamiesShrimali and Lopez-Pamies (2023a) employed it to explain the “pure-shear” fracture test. In two ensuing companion papersShrimali and Lopez-Pamies (2023b, c), the same authors made use of (2) to explain two other types of prominent tests used to study fracture in viscoelastic elastomers: the delayed fracture test and the trousers fracture test.
@@ 1.1 A comment on previous works on the single edge notch fracture test
The first experimental results reported in the literature of single edge notch fracture tests applied to elastomers date back to the 1930s. They correspond to a series of tests carried out by BusseBusse (1934) with specimens made of natural rubber that featured a wide range of sizes for the pre-existing crack.
The early work by BusseBusse (1934) went unnoticed in the mechanics community until Rivlin and ThomasRivlin and Thomas (1953) presented their own single edge notch fracture tests alongside a first theoretical analysis in the 1950s. Their analysis amounts to the approximate evaluation of the criticality condition (1) making use not of local fields but of global quantities — in particular, the separation $l(t)$ between the grips and the corresponding force $P(t)$ at the grips — and of two simplifying constitutive and geometric assumptions:
@@
$-\dfrac{\partial\mathcal{W}}{\partial\mathrm{\Gamma}_{0}}=2AK(\mathrm{\Lambda})W_{{\rm un}}(\mathrm{\Lambda})=T_{c},$ (4)
@@
In 1963, a decade after Rivlin and ThomasRivlin and Thomas (1953) introduced the result (4), GreensmithGreensmith (1963) carried out single edge notch fracture tests on four different types of natural rubbers aimed at determining the numerical factor $K(\mathrm{\Lambda})$ in (4) directly from experimental measurements. The fitting of (4) to his data indicated that, irrespective of the type of rubber that he tested, $K(\mathrm{\Lambda})$ decreases nonlinearly from a value of about $3$ at small stretches when $\mathrm{\Lambda}\approx 1$ to a value somewhat below $2$ at $\mathrm{\Lambda}\approx 3$. Later, in 1972, guided by the experiments of GreensmithGreensmith (1963) and by his own 2D plane stress finite element (FE) simulations of the test, LindleyLindley (1972) proposed the approximation
@@
By now, experiments beyond those of GreensmithGreensmith (1963) are abundant [^4] The single edge notch fracture test has also been of recent popular use to study fracture in elastomers containing a liquid solventKwon et al. (2011); Guo et al. (2016); Lin et al. (2022). in the literatureHamed and Park (1999); Cristiano et al. (2011); Fan et al. (2017); Roucou et al. (2018); Slootman et al. (2020); Menn (2022); Wang and Fan (2023); Xue et al. (2024). However, most of these experiments have been carried out at a fixed constant stretch rate — wherein a global stretch of the form $\mathrm{\Lambda}=1+\dot{\mathrm{\Lambda}}_{0}t$, with fixed $\dot{\mathrm{\Lambda}}_{0}$, is prescribed — which conceals the viscous dissipation that inevitably takes place in these tests. Furthermore, many of these experiments have been carried out with specimens for which the size of the pre-existing crack is comparable to both the specimen width and the specimen thickness, that is, $A\sim H$ and $A\sim B$, so that boundary effects and the 3D character of the mechanical fields are not expected to be negligible.
Beyond the simulations of LindleyLindley (1972), a plurality of FE simulations are also now available in the literatureYeoh (2002); Chen et al. (2017); Kang et al. (2020); Liu et al. (2020). Invariably, all of these simulations assume that the mechanical behavior of elastomers is purely elastic. What is more, they are restricted to 2D plane stress simulations.
@@ 2.1 Initial configuration and kinematics
$\overline{\mathrm{\Omega}}_{0}=\{{\bf X}:{\bf X}\in\mathcal{P}_{0}\setminus\mathrm{\Gamma}_{0}\},$
@@
${\bf F}({\bf X},t)=\nabla{\bf y}({\bf X},t)=\frac{\partial{\bf y}}{\partial{\bf X}}({\bf X},t)$
@@
$\textbf{V}({\bf X},t)=\dot{{\bf y}}({\bf X},t)=\frac{\partial{\bf y}}{\partial t}({\bf X},t).$
@@ 2.2 Constitutive behavior of the elastomer
The specimens are taken to be made of an isotropic incompressible elastomer. Making use of the two-potential formalism Kumar and Lopez-Pamies (2016), we describe its constitutive behavior by two thermodynamic potentials, a free energy density, or stored-energy function, of the form
$\psi({\bf F},{\bf C}^{v})=\left\{\begin{array}[]{ll}\psi^{{\rm Eq}}(I_{1})+\psi^{{\rm NEq}}(I_{1}^{e})&\,{\rm if}\,\,J=1\\ \\ +\infty&\,{\rm else}\end{array}\right.$ (6)
@@
$\phi({\bf F},{\bf C}^{v},\dot{{\bf C}}^{v})=\left\{\begin{array}[]{ll}\dfrac{1}{2}\dot{{\bf C}}^{v}\cdot\mbox{\boldmath$\mathcal{A}$}({\bf F},{\bf C}^{v})\dot{{\bf C}}^{v}&\,{\rm if}\,\,J^{v}=1\\ \\ +\infty&\,{\rm else}\end{array}\right.$ (7)
@@
$I_{1}={\rm tr}\,{\bf C},\quad J=\sqrt{\det{\bf C}},\quad I_{1}^{e}={\rm tr}({\bf C}{{\bf C}^{v}}^{-1}),\quad J^{v}=\sqrt{\det{\bf C}^{v}},$
@@
Granted the two thermodynamic potentials (6) and (7), it follows that the first Piola-Kirchhoff stress tensor ${\bf S}$ at any material point ${\bf X}\in\mathrm{\Omega}_{0}$ and time $t\in[0,T]$ is given by the relation Kumar and Lopez-Pamies (2016)
${\bf S}({\bf X},t)=\frac{\partial\psi}{\partial{\bf F}}({\bf F},{\bf C}^{v}),$
@@
$\left\{\begin{array}[]{l}\dfrac{\partial\psi}{\partial{\bf C}^{v}}({\bf F},{\bf C}^{v})+\dfrac{\partial\phi}{\partial\dot{{\bf C}}^{v}}({\bf F},{\bf C}^{v},\dot{{\bf C}}^{v})={\bf 0}\\ \\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right..$
@@
$\left\{\begin{array}[]{l}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{2\psi^{{\rm NEq}}_{I^{e}_{1}}}{\eta({\bf C},{\bf C}^{v})}\left({\bf C}-\dfrac{1}{3}\left({\bf C}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right)\\ \\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right..$ (9)
@@
For a detailed account of the constitutive relation (8)-(9) and its physical interpretation, the interested reader is referred to Kumar and Lopez-PamiesKumar and Lopez-Pamies (2016). Here, it suffices to make the following remark.
Remark 1
The constitutive relation (8)-(9) is nothing more than a generalization of the classical Zener or standard solid model Zener (1948) to the setting of finite deformations. To see this, note that the rheological representation of the constitutive relation (8)-(9) is given by the Zener rheological representation shown in Fig. 2. Note as well that in the limit of small deformations, as ${\bf F}\rightarrow{\bf I}$, the constitutive relation (8)-(9) reduces to the classical linear viscoelastic Zener model
${\bf S}({\bf X},t)=2\,\mu^{{\rm Eq}}\,\boldsymbol{\varepsilon}+2\,\mu^{{\rm NEq}}(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v})-p{\bf I},$
@@
$\left\{\begin{array}[]{l}\dot{\boldsymbol{\varepsilon}}^{v}({\bf X},t)=\dfrac{\mu^{{\rm NEq}}}{\eta_{0}}\left(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v}-\dfrac{1}{3}{\rm tr}(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v}){\bf I}\right)\\ \\ \boldsymbol{\varepsilon}^{v}({\bf X},0)=\textbf{0}\end{array}\right..$
@@
$\left\{\begin{array}[]{l}\psi^{{\rm Eq}}(I_{1})=\displaystyle\sum\limits_{r=1}^{2}\dfrac{3^{1-\alpha_{r}}}{2\alpha_{r}}\mu_{r}\left(\,I_{1}^{\alpha_{r}}-3^{\alpha_{r}}\right)\vspace{0.2cm}\\ \psi^{{\rm NEq}}(I^{e}_{1})=\displaystyle\sum\limits_{r=1}^{2}\dfrac{3^{1-\beta_{r}}}{2\beta_{r}}\nu_{r}\left(\,{I^{e}_{1}}^{\beta_{r}}-3^{\beta_{r}}\right)\vspace{0.2cm}\\ \eta({\bf C},{\bf C}^{v})=\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})\end{array}\right.$ (10)
@@
$\left\{\begin{array}[]{l}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{\sum\limits_{r=1}^{N}3^{1-\beta_{r}}\nu_{r}{I^{e}_{1}}^{\beta_{r}-1}}{\eta({\bf C},{\bf C}^{v})}\left({\bf C}-\dfrac{1}{3}\left({\bf C}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right)\\ \\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right..$ (12)

### v4

@@ Abstract
Rubber, Elastomers, Adhesives, Cavitation, Fracture
@@ 1 Introduction
As a direct consequence of two elementary observations, [SLP23a] have shown that the original form [RT53, Greensmith55]
$\dfrac{\partial\mathcal{W}}{\partial\mathrm{\Gamma}_{0}}=T_{c}$ (1)
@@
$\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial\mathrm{\Gamma}_{0}}=G_{c}.$ (2)
In expression (1), the left-hand side $\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ denotes the change [^2] The derivative $\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ is to be carried out under conditions of fixed deformation of those parts of the boundary that are not traction-free so that no work is done by the external forces. in total deformation — stored and dissipated [^3] The stored part of the energy is comprised of an equilibrium part $\mathcal{W}^{{\rm Eq}}$ and a non-equilibrium part $\mathcal{W}^{{\rm NEq}}$, the latter representing the part of the stored energy that gets dissipated via viscous deformation eventually. For example, for elastomers whose viscoelastic behavior can be described by the Zener model depicted in Fig. 2, in terms of an equilibrium stored-energy function $\psi^{{\rm Eq}}$, a non-equilibrium stored-energy function $\psi^{{\rm NEq}}$, and a dissipation potential $\phi$, the three parts in the energy decomposition (3) are given by the integrals $\mathcal{W}^{{\rm Eq}}=\displaystyle\int_{\mathrm{\Omega}_{0}}\psi^{{\rm Eq}}\,{\rm d}{\bf X},\qquad\mathcal{W}^{{\rm NEq}}=\displaystyle\int_{\mathrm{\Omega}_{0}}\psi^{{\rm NEq}}\,{\rm d}{\bf X},\qquad\mathcal{W}^{v}=\displaystyle\int_{\mathrm{\Omega}_{0}}\displaystyle\int_{0}^{t}2\phi\,{\rm d}\tau\,{\rm d}{\bf X}$ over the volume $\mathrm{\Omega}_{0}$ occupied by the elastomer in its undeformed configuration to be evaluated at the fields that satisfy the corresponding initial-boundary-value problem; Section 3 describes these calculations in detail for the case of the single edge notch fracture test. — energy
$\mathcal{W}=\underbrace{\mathcal{W}^{{\rm Eq}}+\mathcal{W}^{{\rm NEq}}}_{\text{stored}}+\underbrace{\mathcal{W}^{v}}_{\text{dissipated}}$ (3)
in the specimen at hand with respect to an added surface area ${\rm d}\mathrm{\Gamma}_{0}$ to the pre-existing crack $\mathrm{\Gamma}_{0}$ in its reference state, while the right-hand side $T_{c}$ stands for the critical tearing energy, a characteristic property of the elastomer that depends on the loading history. Traditionally, the critical tearing energy has been written in the form [GentSchultz72, Gent96]
$T_{c}=G_{c}(1+f_{c}),$ (4)
@@
In expression (2), on the other hand, the left-hand side $\partial\mathcal{W}^{{\rm Eq}}/$ $\partial\mathrm{\Gamma}_{0}$ denotes the change not in $\mathcal{W}$ but only in equilibrium elastic energy $\mathcal{W}^{{\rm Eq}}$ stored in the specimen with respect to an added surface area to the pre-existing crack $\mathrm{\Gamma}_{0}$, while the right-hand side involves not $T_{c}$ but only the intrinsic fracture energy $G_{c}$ of the elastomer.
@@
where the right-hand side is to be evaluated at the instance in time at which (2) is attained along the loading path of interest. This last relation clarifies how $T_{c}$ describes the total energy (per unit undeformed fracture area) expended in the tearing process, as it makes explicit the contributions by the creation of new surface ($G_{c}$) and by viscous effects ($\partial\mathcal{W}^{{\rm NEq}}/\partial\mathrm{\Gamma}_{0}-\partial\mathcal{W}^{v}/\partial\mathrm{\Gamma}_{0}$).
@@
Figure 1: Schematic of the single edge notch fracture test for a viscoelastic elastomer illustrating the geometry of the specimen, the applied time-dependent separation $l(t)$ between the grips, and the corresponding force $P(t)$ at the grips. It is standard practice to use the normalized quantities $\mathrm{\Lambda}=l(t)/L$ and $S=P(t)/(HB)$. We refer to them, respectively, as the global stretch and the global nominal stress. The focus of this work is on initial sizes $A$ of the crack (the “notch”) that are larger than the initial thickness $B$ of the specimen.
As a first application of the criticality condition (2), [SLP23a] employed it to explain the “pure-shear” fracture test. In two ensuing companion papers [SLP23b, SLP23c], the same authors made use of (2) to explain two other types of prominent tests used to study fracture in viscoelastic elastomers: the delayed fracture test and the trousers fracture test. More recently, the criticality condition (2) has been recast into a regularized phase-field formulation [KBLP25] and shown to explain additional experimental observations.
@@ 1.1 A comment on previous works on the single edge notch fracture test
The first experimental results reported in the literature of single edge notch fracture tests applied to elastomers date back to the 1930s. They correspond to a series of tests carried out by [Busse34] with specimens made of natural rubber that featured a wide range of sizes for the pre-existing crack.
The early work by [Busse34] went unnoticed in the mechanics community until [RT53] presented their own single edge notch fracture tests alongside a first theoretical analysis in the 1950s. Their analysis amounts to the approximate evaluation of the criticality condition (1) making use not of local fields but of global quantities — in particular, the separation $l(t)$ between the grips and the corresponding force $P(t)$ at the grips — and of two simplifying constitutive and geometric assumptions:
@@
$\dfrac{\partial\mathcal{W}}{\partial\mathrm{\Gamma}_{0}}=2AK(\mathrm{\Lambda})W_{{\rm un}}(\mathrm{\Lambda})=T_{c},$ (6)
@@
In 1963, a decade after [RT53] introduced the result (6), [Greensmith63] carried out single edge notch fracture tests on four different types of natural rubbers aimed at determining the numerical factor $K(\mathrm{\Lambda})$ in (6) directly from experimental measurements. The fitting of (6) to his data indicated that, irrespective of the type of rubber that he tested, $K(\mathrm{\Lambda})$ decreases nonlinearly from a value of about $3$ at small stretches when $\mathrm{\Lambda}\approx 1$ to a value somewhat below $2$ at $\mathrm{\Lambda}\approx 3$. Later, guided by the experiments of [Greensmith63] and by his own 2D plane stress finite element (FE) simulations of the test, [Lindley72] proposed the approximation
@@
By now, experiments beyond those of [Greensmith63] are abundant [^4] The single edge notch fracture test has also been of recent popular use to study fracture in elastomers containing a liquid solvent [Kwon11, Marcellan16, Zhao22]. in the literature [Hamed99, Creton11, Cai17, Mbiakop18, Creton20, LeMenn22, Wang23, Dufresne24]. However, most of these experiments have been carried out at a fixed constant stretch rate — wherein a global stretch of the form $\mathrm{\Lambda}=1+\dot{\mathrm{\Lambda}}_{0}t$, with fixed $\dot{\mathrm{\Lambda}}_{0}$, is prescribed — which conceals the viscous dissipation that inevitably takes place in these tests. Furthermore, many of these experiments have been carried out with specimens for which the size of the pre-existing crack is comparable to both the specimen width and the specimen thickness, that is, $A\sim H$ and $A\sim B$, so that boundary effects and the 3D character of the mechanical fields are not expected to be negligible.
Beyond the simulations of [Lindley72], a plurality of FE simulations are also now available in the literature [Yeoh02, Chen17, Kangetal2020, Liuetal2020]. Invariably, all of these simulations assume that the mechanical behavior of elastomers is purely elastic. What is more, they are restricted to 2D plane stress simulations.
@@
1.2 Clarification of two misconceptions of the criticality condition (2)
Before proceeding with the proposed analysis of the single edge notch fracture test, it is appropriate to close this Introduction by clarifying two misconceptions regarding the criticality condition (2) that have recently appeared in the literature.
The first of these misconceptions is that the equilibrium energy $\mathcal{W}^{{\rm Eq}}$ in the criticality condition (2) is notthe elastic energy stored in the system at equilibrium i.e. the elastic energy after keeping the system which prevail at time $t$ for an infinite long time with fixed boundary conditions” as stated, for instance, in [Persson2024]. Instead, $\mathcal{W}^{{\rm Eq}}$ is the integral over the volume occupied by the elastomer in its undeformed configuration evaluated at the fields that satisfy the corresponding initial-boundary-value problem at the current time $t$. For example, as already noted above, for elastomers whose viscoelastic behavior is described by the Zener model depicted in Fig. 2
$\mathcal{W}^{{\rm Eq}}=\displaystyle\int_{\mathrm{\Omega}_{0}}\psi^{{\rm Eq}}(I_{1})\,{\rm d}{\bf X}\quad{\rm with}\quad I_{1}={\rm tr}\left(\nabla{\bf y}^{T}({\bf X},t)\nabla{\bf y}({\bf X},t)\right),$
where the deformation field ${\bf y}({\bf X},t)$ is solution of the equations of equilibrium at the current time $t$. By definition, the equilibrium energy $\mathcal{W}^{{\rm Eq}}$ depends thus on the entire viscoelastic behavior of the elastomer — implicitly via the deformation field ${\bf y}({\bf X},t)$ — and not just on its behavior at equilibrium. This implies, in particular, that $\mathcal{W}^{{\rm Eq}}$ and hence its derivative $\partial\mathcal{W}^{{\rm Eq}}/\partial\Gamma_{0}$ in the criticality condition (2) can be rate dependent.
The second misconception is that the criticality condition (2) does notonly address the onset of crack propagation” as also stated in [Persson2024]. Instead, the criterion describes the growth of cracks in a complete manner, that is, it describes the growth of a large crack in a viscoelastic elastomer that is subjected to any loading condition, absent inertia. As such, it describes for instance the (uniform or not) speed at which a crack grows or arrests. These kind of predictions have been recently reported in [KBLP25], where the criterion has been recast into a regularized phase-field formulation that is amenable to computations for arbitrary loading conditions.
@@ 2.1 Initial configuration and kinematics
$\overline{\mathrm{\Omega}}_{0}=\{{\bf X}:{\bf X}\in\mathcal{P}_{0}\setminus\mathrm{\Gamma}_{0}\},$ (8)
@@
${\bf F}({\bf X},t)=\nabla{\bf y}({\bf X},t)=\frac{\partial{\bf y}}{\partial{\bf X}}({\bf X},t)$ (12)
@@
$\textbf{V}({\bf X},t)=\dot{{\bf y}}({\bf X},t)=\frac{\partial{\bf y}}{\partial t}({\bf X},t).$ (13)
@@ 2.2 Constitutive behavior of the elastomer
The specimens are taken to be made of an isotropic incompressible elastomer. Making use of the two-potential formalism [KLP16], we describe its constitutive behavior by two thermodynamic potentials, a free energy density, or stored-energy function, of the form
$\psi({\bf F},{\bf C}^{v})=\left\{\begin{array}[]{ll}\hskip-2.84544pt\psi^{{\rm Eq}}(I_{1})+\psi^{{\rm NEq}}(I_{1}^{e})&\,{\rm if}\,\,J=1\\ \hskip-2.84544pt+\infty&\,{\rm else}\end{array}\right$. (14)
@@
$\phi({\bf F},{\bf C}^{v},\dot{{\bf C}}^{v})=\left\{\begin{array}[]{ll}\hskip-2.84544pt\dfrac{1}{2}\dot{{\bf C}}^{v}\cdot\mbox{$\mathcal{A}$}({\bf F},{\bf C}^{v})\dot{{\bf C}}^{v}&\,{\rm if}\,\,J^{v}=1\\ \hskip-2.84544pt+\infty&\,{\rm else}\end{array}\right$. (15)
@@
$I_{1}={\rm tr}\,{\bf C},\quad J=\sqrt{\det{\bf C}},\quad I_{1}^{e}={\rm tr}({\bf C}{{\bf C}^{v}}^{-1}),\quad J^{v}=\sqrt{\det{\bf C}^{v}},$ (16)
@@
Granted the two thermodynamic potentials (14) and (15), it follows that the first Piola-Kirchhoff stress tensor $\bf S$} at any material point ${\bf X}\in\mathrm{\Omega}_{0}$ and time $t\in[0,T]$ is given by the relation [KLP16]
${\bf S}({\bf X},t)=\frac{\partial\psi}{\partial{\bf F}}({\bf F},{\bf C}^{v}),$ (17)
@@
$\left\{\begin{array}[]{l}\dfrac{\partial\psi}{\partial{\bf C}^{v}}({\bf F},{\bf C}^{v})+\dfrac{\partial\phi}{\partial\dot{{\bf C}}^{v}}({\bf F},{\bf C}^{v},\dot{{\bf C}}^{v})={\bf 0}\\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right.$. (18)
@@
$\left\{\hskip-2.84544pt\begin{array}[]{l}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{2\psi^{{\rm NEq}}_{I^{e}_{1}}}{\eta({\bf C},{\bf C}^{v})}\left({\bf C}-\dfrac{1}{3}\left({\bf C}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right)\\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right.\hskip-2.84544pt$. (20)
@@
For a detailed account of the constitutive relation (19)-(20) and its physical interpretation, the interested reader is referred to [KLP16]. Here, it suffices to make the following remark.
Remark 2.1.
The constitutive relation (19)-(20) is nothing more than a generalization of the classical Zener or standard solid model [Zener48] to the setting of finite deformations. To see this, note that the rheological representation of the constitutive relation (19)-(20) is given by the Zener rheological representation shown in Fig. 2. Note as well that in the limit of small deformations, as ${\bf F}\rightarrow{\bf I}$, the constitutive relation (19)-(20) reduces to the classical linear viscoelastic Zener model
${\bf S}({\bf X},t)=2\,\mu^{{\rm Eq}}\,\boldsymbol{\varepsilon}+2\,\mu^{{\rm NEq}}(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v})-p{\bf I},$ (21)
@@
$\left\{\begin{array}[]{l}\dot{\boldsymbol{\varepsilon}}^{v}({\bf X},t)=\dfrac{\mu^{{\rm NEq}}}{\eta_{0}}\left(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v}-\dfrac{1}{3}{\rm tr}(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v}){\bf I}\right)\\ \boldsymbol{\varepsilon}^{v}({\bf X},0)={\bf 0}\end{array}\right.$. (22)
@@
$\left\{\begin{array}[]{l}\psi^{{\rm Eq}}(I_{1})=\displaystyle\sum\limits_{r=1}^{2}\dfrac{3^{1-\alpha_{r}}}{2\alpha_{r}}\mu_{r}\left(\,I_{1}^{\alpha_{r}}-3^{\alpha_{r}}\right)\vskip 5.69046pt\\ \psi^{{\rm NEq}}(I^{e}_{1})=\displaystyle\sum\limits_{r=1}^{2}\dfrac{3^{1-\beta_{r}}}{2\beta_{r}}\nu_{r}\left(\,{I^{e}_{1}}^{\beta_{r}}-3^{\beta_{r}}\right)\vskip 5.69046pt\\ \eta({\bf C},{\bf C}^{v})=\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})\end{array}\right$. (23)
@@
$\left\{\begin{array}[]{l}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{\sum\limits_{r=1}^{N}3^{1-\beta_{r}}\nu_{r}{I^{e}_{1}}^{\beta_{r}-1}}{\eta({\bf C},{\bf C}^{v})}\left({\bf C}-\dfrac{1}{3}\left({\bf C}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right)\\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right.$. (25)